OpenStudy (anonymous):

A system of equations is shown below: n = 3m + 5 n − 2m = 3 What is the solution, in the form (m, n), to the system of equations? (2, 11) (1, 8) (−2, −1) (−3, −4)

OpenStudy (godlovesme):

n = 3m+5 so plug '3m+5' for n in the second formula like (3m-5) -2m = 3

OpenStudy (anonymous):

so would it be c? :E

OpenStudy (godlovesme):

one sec

OpenStudy (anonymous):

ok

OpenStudy (godlovesme):

how did u get C? i got different answer O-o

OpenStudy (anonymous):

well i did it wrong, i think i used elimination (i suck at math) but what did you get, could you explain it?(:

OpenStudy (godlovesme):

yes i will :D (3m-5)-2m = 3 did u understand what i did ^^ ?

OpenStudy (anonymous):

yes

OpenStudy (godlovesme):

so (3m-5)-2m= 3 3m-2m -5 = 3 1m -5 + 5 = 3+5 m= 8

OpenStudy (anonymous):

wait, so it would be B (1,8)

OpenStudy (godlovesme):

hold on i think i did something wrong o-o

OpenStudy (anonymous):

ok hahah

OpenStudy (godlovesme):

wait it was +5 NOT -5 my bad :/

OpenStudy (anonymous):

oh its okay

OpenStudy (godlovesme):

ok here (3m + 5) -2m = 3 3m -2m +5 = 3 1m +5-5= 2-5 m= -2

OpenStudy (anonymous):

okay so the answer is c?

OpenStudy (godlovesme):

yes no need to solve for n ;) :P

OpenStudy (anonymous):

thanks! C it is!

OpenStudy (godlovesme):

you're welcome :)